# Measure theoretic properties of large products of consecutive partial quotients

### v1

@@ 4.1. Convergence case
Take $n\in\mathbb{N}_{\geq\ell}$. We estimate $\mathfrak{m}(A_{n,k})$ by considering two cases according to whether the restricted blocks overlap or not.
@@
    $\mathfrak{m}(A_{n,k})\asymp_{\ell}\frac{\log^{\ell+k-n-1}\varphi(n)}{\varphi(n)}.$
@@
    $\mathfrak{m}(A_{n,k})\asymp_{\ell}\sum_{\begin{subarray}{c}b_{k}\cdots b_{k+\ell-1}\geq\varphi(n)\\ b_{n}\cdots b_{n+\ell-1}\geq\varphi(n)\end{subarray}}\frac{1}{b_{k}^{2}\cdots b_{k+\ell-1}^{2}b_{k+\ell}^{2}\cdots b_{n+\ell-1}^{2}}\asymp\frac{\log^{j-1}\varphi(n)}{\varphi(n)}.$
Proof of Theorem 1.5. Convergence case..
@@
$\asymp(n-\ell)\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}+\frac{1}{\varphi(n)}\sum_{k=n-\ell+1}^{n-1}\log^{k+\ell-n-1}\varphi(n)$
$\asymp\frac{n\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}+\frac{\log^{\ell-2}\varphi(n)}{\varphi(n)}.$
@@ 4.2. Divergence case
If this was not the case, then every irrational number with unbounded partial quotients would belong to $\mathcal{F}_{\ell}(\varphi)$ and, thus, $\mathfrak{m}(\mathcal{F}_{\ell}(\varphi))=1$ by the Borel-Bernstein Theorem. Second, we may assume that  [^1] There seems to be a gap in the proof of this assertion for $\ell=2$ in [26]. Consider $10<\alpha<$ 1 and $\varphi(n)=n(\log n)^{\alpha}$. Then, $\sum_{n}\varphi(n)^{-1}=\infty$ and $\varphi(n)<n\log\varphi(n)$ for all large $n$. According to [26, Section 3.2], we should consider the function $\psi$ given by $\psi(n)=\max\{\varphi(n),n\log\varphi(n)\}$. However, for all large $n$ we have $n\log\psi(n)=n\log(n\log\varphi(n))>n\log\varphi(n)=\psi(n)$. This minor gap does not affect validity of that paper at all.
(13) $\varphi(n)\geq n\log^{\ell-1}\varphi(n)\quad\text{ for all }n\in\mathbb{N}.$
@@
We can trivially modify at most finitely many terms of $\psi$ to ensure that it is non-decreasing. Furthermore, since $\psi\geq\varphi$, we have $\mathcal{E}_{\ell}(\psi)\subseteq\mathcal{E}_{\ell}(\varphi)$ and $\psi(n)\to\infty$ as $n\to\infty$. It is also clear that $\psi(n)\geq n\log^{\ell-1}\psi(n)$ for all $n\in\mathbb{N}$, because $x\mapsto\frac{x}{\log^{\ell-1}x}$ is strictly increasing for large $x$. Lastly, let us show that
$\sum_{n=1}^{\infty}n\left(\frac{\log^{\ell-1}\psi(n)}{\psi(n)}\right)^{2}+\frac{\log^{\ell-2}\psi(n)}{\psi(n)}=\infty.$
@@
$\sum_{n=1}^{\infty}n\left(\frac{\log^{\ell-1}\psi(n)}{\psi(n)}\right)^{2}+\frac{\log^{\ell-2}\psi(n)}{\psi(n)}$ $\geq\sum_{k=1}^{\infty}\sum_{\left[\frac{m_{k+1}}{2}\right]\leq n\leq m_{k+1}}n\left(\frac{\log^{\ell-1}\psi(n)}{\psi(n)}\right)^{2}$
$\geq\sum_{k=1}^{\infty}\frac{m_{k+1}}{2}\sum_{\left[\frac{m_{k+1}}{2}\right]\leq n\leq m_{k+1}}\left(\frac{\log^{\ell-1}\psi(n)}{\psi(n)}\right)^{2}$
$\geq\sum_{k=1}^{\infty}\frac{m_{k+1}}{2}\sum_{\left[\frac{m_{k+1}}{2}\right]\leq n\leq m_{k+1}}\left(\frac{\log^{\ell-1}x_{m_{k+1}}}{x_{m_{k+1}}}\right)^{2}$
@@
Take $t\in\mathbb{N}$, $\mathbf{b}=(b_{1},\ldots,b_{t})\in\mathbb{N}^{t}$, write $I_{t}=I_{t}(\mathbf{b})$, and let $M=M(\mathbf{b})\in\mathbb{N}_{\geq t+1}$ be such that $\varphi(M)>\left(\max\{b_{1},\ldots,b_{t}\}\right)^{\ell}$.
@@
$\mathfrak{m}\left(A_{n}\cap I_{t}\right)\gg_{\ell}\mathfrak{m}(I_{t})\left(\frac{n\log^{2(\ell-1)}\varphi(n)}{\varphi(n)^{2}}+\frac{\log^{\ell-2}\varphi(n)}{\varphi(n)}\right).$
@@
$B_{n}:=\left\{x\in I_{t}:\begin{matrix}a_{n-1}(x)\cdots a_{n+\ell-2}(x)\geq\varphi(n),\\ a_{n}(x)\cdots a_{n+\ell-1}(x)\geq\varphi(n)\end{matrix}\right\}.$
For each $i\in\mathbb{N}$ such that $M\leq i<n-1$ and $i\equiv n-1\pmod{\ell}$, let $B_{i}$ be the set of those $x\in I_{t}$ satisfying the next properties:
@@
    For each $k\in\mathbb{N}$ such that $0<k<\frac{n-1-i}{\ell}$ we have $a_{i+k\ell}(x)\cdots a_{i+(k+1)\ell-1}(x)<\varphi(n)$.
@@
    $a_{n-1}(x)\cdots a_{n+\ell-2}(x)<\varphi(n)$ and $a_{n}(x)\cdots a_{n+\ell-1}(x)\geq\varphi(n)$.
Note that we are imposing restrictions on non-overlapping contiguous blocks of length $\ell$. Clearly, we have
(14) $B_{n}\cup\bigcup_{\begin{subarray}{c}M\leq i<n-1\\ i\equiv n-1(\ell)\end{subarray}}B_{i}\subseteq A_{n}\cap I_{t},$
@@
(15) $\mathfrak{m}(B_{n})\asymp_{\ell}\frac{\log^{\ell-1}\varphi(n)}{\varphi(n)}\mathfrak{m}(I_{t}).$
Take $i$ such that $i\equiv n-1\pmod{\ell}$ and $M\leq i<n-1$. As in the convergence case, in order to estimate $\mathfrak{m}(B_{i})$ it suffices to estimate the series
$\sum_{a_{i}\cdots a_{i+\ell-1}\geq\varphi(n)}\,\sum_{\begin{subarray}{c}a_{i+k\ell}\cdots a_{i+(k+1)\ell-1}<\varphi(n)\\ 0<k<\frac{n-1-i}{\ell}\end{subarray}}\,\sum_{\begin{subarray}{c}a_{n-1}\cdots a_{n+\ell-2}<\varphi(n)\\ a_{n}\cdots a_{n+\ell-1}\geq\varphi(n)\end{subarray}}\frac{1}{a_{i}^{2}\cdots a_{n+\ell-2}^{2}}.$
@@
$\sum_{a_{i}\cdots a_{i+\ell-1}\geq\varphi(n)}\frac{1}{a_{i}^{2}\cdots a_{i+\ell-1}^{2}}\;\sum_{\begin{subarray}{c}a_{n-1}\cdots a_{n+\ell-2}<\varphi(n)\\ a_{n}\cdots a_{n+\ell-1}\geq\varphi(n)\end{subarray}}\frac{1}{a_{n-1}^{2}\cdots a_{n+\ell+1}^{2}}\asymp\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}.$
Also, for every integer $k$ with $0<k<\frac{n-1-i}{\ell}$ we have
$\sum_{a_{i+k\ell}\cdots a_{i+(k+1)\ell-1}<\varphi(n)}\frac{1}{a_{i+k\ell}^{2}\cdots a_{i+(k+1)\ell-1}^{2}}\gg 1-\frac{\log^{\ell-1}\varphi(n)}{2\varphi(n)}\geq 1-\frac{1}{2n}.$
@@
$\mathfrak{m}(B_{i})\gg\mathfrak{m}(I_{t})\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi(n)}\left(1-\frac{1}{2n}\right)^{n-i}$
@@
(16) $\mathfrak{m}\left(\bigcup_{\begin{subarray}{c}M\leq i<n-1\\ i\equiv n-1(\ell)\end{subarray}}B_{i}\right)\gg\mathfrak{m}(I_{t})\frac{n\log^{2(\ell-1)}\varphi(n)}{\varphi(n)}.$
@@
(17) $\sum_{M\leq m<n\leq N}\mathfrak{m}\left(A_{m}\cap A_{n}\cap I_{t}\right)\ll_{\ell}\sum_{n=M}^{N}\mathfrak{m}\left(A_{n}\cap I_{t}\right)+\mathfrak{m}(I_{t})^{-1}\left(\sum_{n=M}^{N}\mathfrak{m}\left(A_{n}\cap I_{t}\right)\right)^{2}.$
@@
$\mathcal{I}_{1}:=\{(m,n)\in\mathcal{I}:M\leq m\leq n-(2\ell-1)\},\quad\mathcal{I}_{2}:=\{(m,n)\in\mathcal{I}:n-2\ell+2\leq m\leq n-1\}\}$
@@
$S_{2}\leq 2\ell\sum_{n=1}^{N}\mathfrak{m}\left(A_{n}\cap I_{t}\right).$
@@
    • $a_{m}(x)\cdots a_{m+\ell-1}(x)\geq\varphi(m)$,
    • For some $k<m$ we have $a_{k}(x)\cdots a_{k+\ell-1}(x)\geq\varphi(n)$,
    • $a_{n}(x)\cdots a_{n+\ell-1}(x)\geq\varphi(n)$.
@@
    • For some $j<m$ we have $a_{j}(x)\cdots a_{j+\ell-1}(x)\geq\varphi(m)$,
    • $a_{m}(x)\cdots a_{m+\ell-1}(x)\geq\varphi(n)$,
    • $a_{n}(x)\cdots a_{n+\ell-1}(x)\geq\varphi(n)$.
@@
    • For some $j<m$ we have $a_{j}(x)\cdots a_{j+\ell-1}\geq\varphi(m)$,
    • $a_{m}(x)\cdots a_{m+\ell-1}(x)\geq\varphi(m)$,
    • For some $m<k<n$ we have $a_{k}(x)\cdots a_{k+\ell-1}\geq\varphi(n)$,
    • $a_{n}(x)\cdots a_{m+\ell-1}(x)\geq\varphi(m)$.
@@
(18) $\mathfrak{m}(U_{1}(m,n))\ll\mathfrak{m}(I_{t})\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)},$
(19) $\mathfrak{m}(U_{2}(m,n))\ll\mathfrak{m}(I_{t})\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)},$
$\mathfrak{m}(U_{3}(m,n))$ $\ll\mathfrak{m}(I_{t})\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}+$
(20) $\quad\;+\mathfrak{m}(I_{t})\left(\frac{n\log\varphi(n)}{\varphi^{2}(n)}+\frac{\log^{\ell-2}\varphi(n)}{\varphi(n)}\right)\left(\frac{m\log\varphi(m)}{\varphi^{2}(m)}+\frac{\log^{\ell-2}\varphi(n)}{\varphi(m)}\right).$
Let us show (18). When $M\leq k\leq m-\ell$, we apply (5) to get
$\sum_{\begin{subarray}{c}a_{k}\cdots a_{k+\ell-1}\geq\varphi(n)\\ a_{m}\cdots a_{m+\ell-1}\geq\varphi(m)\\ a_{n}\cdots a_{n+\ell-1}\geq\varphi(n)\end{subarray}}\frac{1}{a_{k}^{2}\cdots a_{k+\ell-1}^{2}a_{m}^{2}\cdots a_{m+\ell-1}^{2}a_{n}^{2}\cdots a_{n+\ell-1}^{2}}\asymp\frac{\log\varphi(m)}{\varphi(m)}\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}.$
For $m-\ell+1\leq k\leq m-1$ we have
$\sum_{\begin{subarray}{c}a_{k}\cdots a_{k+\ell-1}\geq\varphi(n)\\ a_{m}\cdots a_{m+\ell-1}\geq\varphi(m)\\ a_{n}\cdots a_{n+\ell-1}\geq\varphi(n)\end{subarray}}\frac{1}{a_{k}^{2}\cdots a_{m+\ell-1}^{2}a_{n}^{2}\cdots a_{n+\ell-1}^{2}}$ $\leq\sum_{\begin{subarray}{c}a_{k}\cdots a_{k+\ell-1}\geq\varphi(n)\\ a_{n}\cdots a_{n+\ell-1}\geq\varphi(n)\end{subarray}}\frac{1}{a_{k}^{2}\cdots a_{m+\ell-1}^{2}a_{n}^{2}\cdots a_{n+\ell-1}^{2}}$
$\ll\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}.$

### v2

@@ 4.1. Convergence case
Take $n\in\mathbb{N}_{\geq\ell}$. We show that
(14) $\frac{n\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}+\frac{\log^{2(\ell-2)}\varphi(n)}{\varphi(n)}<\infty\quad\text{ implies }\quad\mathfrak{m}(\mathcal{F}_{\ell}(\varphi))=0.$
By the end of the proof, we say how to modify the argument to obtain the convergence case in Theorem 1.5. We estimate $\mathfrak{m}(A_{n,k})$ by considering two cases according to whether the restricted blocks overlap or not.
@@
    $\mathfrak{m}(A_{n,k})\ll_{\ell}\frac{\log^{2(\ell+k-n-1)}\varphi(n)}{\varphi(n)}.$
@@
    $\mathfrak{m}(A_{n,k})\asymp_{\ell}\sum_{\begin{subarray}{c}b_{k}\cdots b_{k+\ell-1}\geq\varphi(n)\\ b_{n}\cdots b_{n+\ell-1}\geq\varphi(n)\end{subarray}}\frac{1}{b_{k}^{2}\cdots b_{k+\ell-1}^{2}b_{k+\ell}^{2}\cdots b_{n+\ell-1}^{2}}\ll\frac{\log^{2(j-1)}\varphi(n)}{\varphi(n)}.$
Proof of (14) .
@@
$\ll(n-\ell)\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}+\frac{1}{\varphi(n)}\sum_{k=n-\ell+1}^{n-1}\log^{2(k+\ell-n-1)}\varphi(n)$
$\asymp\frac{n\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}+\frac{\log^{2(\ell-2)}\varphi(n)}{\varphi(n)}.$
@@
For the convergence part in Theorem 1.5, set $\ell=3$ and use (6) instead of (7) in the previous argument.
@@ 4.2. Divergence case
If this was not the case, then every irrational number with unbounded partial quotients would belong to $\mathcal{F}_{3}(\varphi)$ and, thus, $\mathfrak{m}(\mathcal{F}_{3}(\varphi))=1$ by the Borel-Bernstein Theorem. Second, we may assume that  [^1] There seems to be a gap in the proof of this assertion for $\ell=2$ in [28]. Consider $10<\alpha<$ 1 and $\varphi(n)=n(\log n)^{\alpha}$. Then, $\sum_{n}\varphi(n)^{-1}=\infty$ and $\varphi(n)<n\log\varphi(n)$ for all large $n$. According to [28, Section 3.2], we should consider the function $\psi$ given by $\psi(n)=\max\{\varphi(n),n\log\varphi(n)\}$. However, for all large $n$ we have $n\log\psi(n)=n\log(n\log\varphi(n))>n\log\varphi(n)=\psi(n)$. This minor gap does not affect validity of that paper at all.
(15) $\varphi(n)\geq n\log^{2}\varphi(n)\quad\text{ for all }n\in\mathbb{N}.$
@@
We can trivially modify at most finitely many terms of $\psi$ to ensure that it is non-decreasing. Furthermore, since $\psi\geq\varphi$, we have $\mathcal{E}_{3}(\psi)\subseteq\mathcal{E}_{3}(\varphi)$ and $\psi(n)\to\infty$ as $n\to\infty$. It is also clear that $\psi(n)\geq n\log^{2}\psi(n)$ for all $n\in\mathbb{N}$, because $x\mapsto\frac{x}{\log^{2}x}$ is strictly increasing for large $x$. Lastly, let us show that
$\sum_{n=1}^{\infty}n\left(\frac{\log^{2}\psi(n)}{\psi(n)}\right)^{2}+\frac{\log\psi(n)}{\psi(n)}=\infty.$
@@
$\sum_{n=1}^{\infty}n\left(\frac{\log^{2}\psi(n)}{\psi(n)}\right)^{2}+\frac{\log\psi(n)}{\psi(n)}$ $\geq\sum_{k=1}^{\infty}\sum_{\left[\frac{m_{k+1}}{2}\right]\leq n\leq m_{k+1}}n\left(\frac{\log^{2}\psi(n)}{\psi(n)}\right)^{2}$
$\geq\sum_{k=1}^{\infty}\frac{m_{k+1}}{2}\sum_{\left[\frac{m_{k+1}}{2}\right]\leq n\leq m_{k+1}}\left(\frac{\log^{2}\psi(n)}{\psi(n)}\right)^{2}$
$\geq\sum_{k=1}^{\infty}\frac{m_{k+1}}{2}\sum_{\left[\frac{m_{k+1}}{2}\right]\leq n\leq m_{k+1}}\left(\frac{\log^{2}x_{m_{k+1}}}{x_{m_{k+1}}}\right)^{2}$
@@
Take $t\in\mathbb{N}$, $\mathbf{b}=(b_{1},\ldots,b_{t})\in\mathbb{N}^{t}$, write $I_{t}=I_{t}(\mathbf{b})$, and let $M=M(\mathbf{b})\in\mathbb{N}_{\geq t+1}$ be such that $\varphi(M)>\left(\max\{b_{1},\ldots,b_{t}\}\right)^{3}$.
@@
$\mathfrak{m}\left(A_{n}\cap I_{t}\right)\gg\mathfrak{m}(I_{t})\left(\frac{n\log^{4}\varphi(n)}{\varphi^{2}(n)}+\frac{\log\varphi(n)}{\varphi(n)}\right).$
@@
$B_{n}:=\left\{x\in I_{t}:\begin{matrix}a_{n-1}(x)a_{n}(x)a_{n+1}(x)\geq\varphi(n),\\ a_{n}(x)a_{n+1}(x)a_{n+2}(x)\geq\varphi(n)\end{matrix}\right\},$
For each $i\in\mathbb{N}$ such that $M\leq i<n-1$ and $i\equiv n-1\pmod{3}$, let $B_{i}$ be the set of those $x\in I_{t}$ satisfying the next properties:
@@
    For each $k\in\mathbb{N}$ such that $0<k<\frac{n-1-i}{3}$ we have $a_{i+3k}(x)a_{i+1+3k}(x)a_{i+2+3k}(x)<\varphi(n)$.
@@
    $a_{n-1}(x)a_{n}(x)a_{n+1}(x)<\varphi(n)$ and $a_{n}(x)a_{n+1}(x)a_{n+2}(x)\geq\varphi(n)$.
Note that we are imposing restrictions on non-overlapping contiguous blocks of length $3$. Clearly, we have
(16) $B_{n}\cup\bigcup_{\begin{subarray}{c}M\leq i<n-1\\ i\equiv n-1(3)\end{subarray}}B_{i}\subseteq A_{n}\cap I_{t},$
@@
(17) $\mathfrak{m}(B_{n})\asymp\frac{\log\varphi(n)}{\varphi(n)}\mathfrak{m}(I_{t}).$
Take $i$ such that $i\equiv n-1\pmod{3}$ and $M\leq i<n-1$. As in the convergence case, in order to estimate $\mathfrak{m}(B_{i})$ it suffices to estimate the series
$\sum_{a_{i}a_{i+1}a_{i+2}\geq\varphi(n)}\,\sum_{\begin{subarray}{c}a_{i+3k}a_{i+1+3k}a_{i+2+3k}<\varphi(n)\\ 0<k<\frac{n-1-i}{3}\end{subarray}}\,\sum_{\begin{subarray}{c}a_{n-1}a_{n}a_{n+1}<\varphi(n)\\ a_{n}a_{n+1}a_{n+2}\geq\varphi(n)\end{subarray}}\frac{1}{a_{i}^{2}a_{i+1}^{2}a_{i+2}^{2}\cdots a_{n-1}^{2}a_{n}^{2}a_{n+1}^{2}}.$
@@
$\sum_{a_{i}a_{i+1}a_{i+2}\geq\varphi(n)}\frac{1}{a_{i}^{2}a_{i+1}^{2}a_{i+2}^{2}}\;\sum_{\begin{subarray}{c}a_{n-1}a_{n}a_{n+1}<\varphi(n)\\ a_{n}a_{n+1}a_{n+2}\geq\varphi(n)\end{subarray}}\frac{1}{a_{n-1}^{2}a_{n}^{2}a_{n+1}^{2}}\asymp\frac{\log^{4}\varphi(n)}{\varphi^{2}(n)}.$
Also, for every integer $k$ with $0<k<\frac{n-1-i}{3}$ we have
$\sum_{a_{i+3k}a_{i+1+3k}a_{i+2+3k}<\varphi(n)}\frac{1}{a_{i+3k}^{2}a_{i+1+3k}^{2}a_{i+2+3k}^{2}}\gg 1-\frac{\log^{2}\varphi(n)}{2\varphi(n)}\geq 1-\frac{1}{2n}.$
@@
$\mathfrak{m}(B_{i})\gg\mathfrak{m}(I_{t})\frac{\log^{4}\varphi(n)}{\varphi(n)}\left(1-\frac{1}{2n}\right)^{n-i}$
@@
(18) $\mathfrak{m}\left(\bigcup_{\begin{subarray}{c}M\leq i<n-1\\ i\equiv n-1(3)\end{subarray}}B_{i}\right)\gg\mathfrak{m}(I_{t})\frac{n\log^{4}\varphi(n)}{\varphi^{2}(n)}.$
@@
(19) $\sum_{M\leq m<n\leq N}\mathfrak{m}\left(A_{m}\cap A_{n}\cap I_{t}\right)\ll\sum_{n=M}^{N}\mathfrak{m}\left(A_{n}\cap I_{t}\right)+\mathfrak{m}(I_{t})^{-1}\left(\sum_{n=M}^{N}\mathfrak{m}\left(A_{n}\cap I_{t}\right)\right)^{2}.$
@@
$\mathcal{I}_{1}:=\{(m,n)\in\mathcal{I}:M\leq m\leq n-5\},\quad\mathcal{I}_{2}:=\{(m,n)\in\mathcal{I}:n-4\leq m\leq n-1\}\}$
@@
$S_{2}\leq 6\sum_{n=M}^{N}\mathfrak{m}\left(A_{n}\cap I_{t}\right).$
@@
    • $a_{m}(x)a_{m+1}(x)a_{m+2}(x)\geq\varphi(m)$,
    • For some $k<m$ we have $a_{k}(x)a_{k+1}(x)a_{k+2}(x)\geq\varphi(n)$,
    • $a_{n}(x)a_{n+1}(x)a_{n+2}(x)\geq\varphi(n)$.
@@
    • For some $j<m$ we have $a_{j}(x)a_{j+1}(x)a_{j+2}(x)\geq\varphi(m)$,
    • $a_{m}(x)a_{m+1}(x)a_{m+2}(x)\geq\varphi(n)$,
    • $a_{n}(x)a_{n+1}(x)a_{n+2}(x)\geq\varphi(n)$.
@@
    • For some $j<m$ we have $a_{j}(x)a_{j+1}(x)a_{j+2}(x)\geq\varphi(m)$,
    • $a_{m}(x)a_{m+1}(x)a_{m+2}(x)\geq\varphi(m)$,
    • For some $m<k<n$ we have $a_{k}(x)a_{k+1}(x)a_{k+2}(x)\geq\varphi(n)$,
    • $a_{n}(x)a_{n+1}(x)a_{n+2}(x)\geq\varphi(n)$.
@@
(20) $\mathfrak{m}(U_{1}(m,n))\ll\mathfrak{m}(I_{t})\frac{\log^{4}\varphi(n)}{\varphi^{2}(n)},$
(21) $\mathfrak{m}(U_{2}(m,n))\ll\mathfrak{m}(I_{t})\frac{\log^{4}\varphi(n)}{\varphi^{2}(n)},$
$\mathfrak{m}(U_{3}(m,n))$ $\ll\mathfrak{m}(I_{t})\frac{\log^{4}\varphi(n)}{\varphi^{2}(n)}+$
(22) $\quad\;+\mathfrak{m}(I_{t})\left(\frac{n\log^{4}\varphi(n)}{\varphi^{2}(n)}+\frac{\log\varphi(n)}{\varphi(n)}\right)\left(\frac{m\log^{4}\varphi(m)}{\varphi^{2}(m)}+\frac{\log\varphi(m)}{\varphi(m)}\right).$
Let us show (20). When $M\leq k\leq m-3$, we apply (5) to get
$\sum_{\begin{subarray}{c}a_{k}a_{k+1}a_{k+2}\geq\varphi(n)\\ a_{m}a_{m+1}a_{m+2}\geq\varphi(m)\\ a_{n}a_{n+1}a_{n+2}\geq\varphi(n)\end{subarray}}\frac{1}{a_{k}^{2}a_{k+1}^{2}a_{k+2}^{2}a_{m}^{2}a_{m+1}^{2}a_{m+2}^{2}a_{n}^{2}a_{n+1}^{2}a_{n+2}^{2}}\asymp\frac{\log\varphi(m)}{\varphi(m)}\frac{\log^{4}\varphi(n)}{\varphi^{2}(n)}.$
For $m-2\leq k\leq m-1$ we have
$\sum_{\begin{subarray}{c}a_{k}a_{k+1}a_{k+2}\geq\varphi(n)\\ a_{m}a_{m+1}a_{m+2}\geq\varphi(m)\\ a_{n}a_{n+1}a_{n+2}\geq\varphi(n)\end{subarray}}\frac{1}{a_{k}^{2}\cdots a_{m+2}^{2}a_{n}^{2}a_{n+1}^{2}a_{n+2}^{2}}$ $\leq\sum_{\begin{subarray}{c}a_{k}a_{k+1}a_{k+2}\geq\varphi(n)\\ a_{n}a_{n+1}a_{n+2}\geq\varphi(n)\end{subarray}}\frac{1}{a_{k}^{2}\cdots a_{m+2}^{2}a_{n}^{2}a_{n+1}^{2}a_{n+2}^{2}}$
$\ll\frac{\log^{4}\varphi(n)}{\varphi^{2}(n)}.$